OpenStudy (anonymous):

Solve for the variable c in this equation. a(b – c) = d Solve for the variable c in this equation. a(b – c) = d

OpenStudy (anonymous):

The formula for the area of a triangle is . What is the length of the base, b, of a triangle with an area, A, of 110 m2 and a height, h, of 10 m? 2200 m 11 m 22 m 1100 m

OpenStudy (anonymous):

i need help on that question to

OpenStudy (anonymous):

First divide a from both sides. then subtract b, now I would multiple both sides by -1 to get a nice positive c on the left side! We end up with \[c=-(\frac{ d-b }{ a})\]

OpenStudy (anonymous):

[distribute the a] ac - ab = d [add ab to both sides] ac = d + ab [divide both sides by a] c = (d + ab)/a can also be written as c = d/a + b

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Triangle problem: remember triangle area is \[ \frac{ Base * Height }{ 2 }\] Let's plug in what we know... \[Area 110m^2 = \frac{ base* 10m }{ 2 }\] Solve it and we get 22m!

OpenStudy (anonymous):

thx